a. Poisson distribution was used to model the number of arrivals per minute at a bank located in the central business district of a city. Suppose that the actual arrivals per minute were observed in 200 01e. minute periods over the course of a week are given below. Determine whether the number of arrivals per minute follows a Poisson distribution Frequency 14 31 47 41 29 21 10 Arrivals 2 6 8 Total 1 -8 CI- b. In the z test, a hypothesized distribution more likely to accepted at a- 005 than at (1-E 200 -001? Explain your answer
Answer #1

Answer:

x

f

f*x

0

14

0

1

31

31

2

47

94

3

41

123

4

29

116

5

21

105

6

10

60

7

5

35

8

2

16

Total

200

580

mean = 580/200=2.9

λ= 2.9

P(X=x) = e λx / x!    for x = 0, 1, ....8

P(X = 0) = 0.0550

POISSON.DIST Probabilities Table

X

P(X)

0

0.0550

1

0.1596

2

0.2314

3

0.2237

4

0.1622

5

0.0940

6

0.0455

7

0.0188

8

0.0068

x

f

f*x

p

200*p= expected value

0

14

0

0.0550

11

1

31

31

0.1596

31.92

2

47

94

0.2314

46.28

3

41

123

0.2237

44.74

4

29

116

0.1622

32.44

5

21

105

0.0940

18.8

6

10

60

0.0455

9.1

7

5

35

0.0188

3.76

8

2

16

0.0068

1.36

Total

200

580

199.4

Ho: number of arrivals follows Poisson distribution

H1: number of arrivals not follows Poisson distribution

(O - E)2

Goodness of Fit Test

observed

expected

O - E

(O - E)² / E

14

11.000

3.000

0.818

31

31.920

-0.920

0.027

47

46.280

0.720

0.011

41

44.740

-3.740

0.313

29

32.440

-3.440

0.365

21

18.800

2.200

0.257

10

9.100

0.900

0.089

5

3.760

1.240

0.409

2

1.360

0.640

0.301

Total

200

199.400

0.600

2.590

Chi square

2.590

DF

8

Calculated chi square = 2.590 is < critical chi square at 0.05 level of significance value 15.51.

Ho is not rejected at 0.05 level.

Therefore Ho is not rejected at 0.01 level also.( critical value for 0.01 level is 20.09)